OpenStudy (anonymous):

Researchers studying fish populations in Sunset Lake caught, marked, and then released 232 bass. Later, a sample of 329 bass were caught and examined. Of these, 16 were found to be marked. Use the proportion below to estimate the total bass population in Sunset Lake. Marked bass in sample/Total bass in sample = Marked bass in lake/ Total bass in lake

OpenStudy (mathmate):

The given formula is only an approximation to the real answer, but here it is anyway: 329/16=232/total => total = 232/16*329=4770.5 say 4771

OpenStudy (anonymous):

@mathmate 329/16 is 20.5625

OpenStudy (anonymous):

this is wrong

OpenStudy (anonymous):

maybe the answer is not but the method is

OpenStudy (anonymous):

Help?????????????????????????????????????????????????????????????????????????????????????????????????????????????

OpenStudy (mathmate):

This is the proportion equation. \( \frac{329}{16}=\frac{total}{232} \) [note correction, my bad] => total = \(\frac{232*329}{16}=4770.5 say 4771 \) [no change] Did you try 4770 or 4771 as an answer

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

thank you for clearing that up

OpenStudy (anonymous):

4770

OpenStudy (anonymous):

where do u get 220 n 222

OpenStudy (anonymous):

no the last one he did

OpenStudy (anonymous):

because the first one was incorrect

hartnn (hartnn):

\(\huge \frac{232*329}{16}=4770.5 say 4771\)

OpenStudy (anonymous):

count six from here and that is where it is

hartnn (hartnn):

\(\huge \frac{329}{16}=\frac{total}{232}\)

OpenStudy (anonymous):

no it says 222 n 220

hartnn (hartnn):

zoom in your web page.

OpenStudy (anonymous):

ok